OpenStudy (anonymous):

Write the equation of the circle with the given center and radius. Center: (0,2); radius: 5

OpenStudy (anonymous):

please walk me through the steps I cant remember how to do it

OpenStudy (anonymous):

x^2 + (y-2)^2 = 5^2

OpenStudy (anonymous):

is that the equation or the answer?

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

okay well thank you very much (:

OpenStudy (anonymous):

Welcome. :-)

OpenStudy (anonymous):

what about this one? center: (-4,-5) ; radius; Square root 2

OpenStudy (anonymous):

@SweetsAndEverything @electrokid

OpenStudy (anonymous):

(x + 4)^2 + (y+5)^2 = 2

OpenStudy (anonymous):

Just memorize the formula and you will know the answer.

OpenStudy (anonymous):

@kaitlin_07 The equation for circle is: \[(x-h)^{2}+(y-k)^{2}=r ^{2}\] It'll be like this: \[[x-(-4)]^{2}+[y-(-5)]^{2}=\sqrt{2}\] \[[x+4)]^{2}+[y+5)]^{2}=2\]